# Non-Adaptive Multi-Stage Algorithm for Group Testing with Prior Statistics

### v2

@@ Abstract
In this paper, we propose an efficient multi-stage algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics. The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes. We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure. Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal Maximum A Posteriori (MAP) performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms. Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
@@ I Introduction
In this work, we introduce a practical non-adaptive Multi-Stage GT (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. Our numerical results demonstrate that in practical regimes for COVID-19 [34, 35] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
The rest of this paper is organized as follows. Section II formally describes the GT model with correlated prior statistics. Section III presents the MSGT algorithm and the analytical results. Section IV details the simulation evaluation. Finally, Section V provides concluding remarks and future directions.
@@ III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the List Viterbi Algorithm (LVA) [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix -A.
@@ III-B Analytical Results
Subsequently, the probability of success satisfies is given by
@@
The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that we conjecture outperforms ML. This can be assumed since LVA used for this step is an optimal MAP estimator [32].
@@
Let $\eta$ denote the average success probability of ML decoder. For GT with correlated prior information and $\gamma\geq 1$, MSGT can achieve an average success probability equal or greater than $\eta$, if
@@
Consider a group test for a population of size $N$ items out of which $K$ are defective and a Bernoulli testing matrix. The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
@@ III-C Discussion
IV Numerical Evaluation
@@
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection IV-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection IV-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [50], which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [34, 35], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 51]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
IV-1 Theoretical Analysis
@@
In Fig. 4(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and on the worst-case that were calculated in Theorems 1 and 2 respectively. Note that the worst-case scenario regarding MSGT is when LVA filters the correct set of defective items. That may happen if the number of possibly defective items exceeds the threshold $\gamma K$. Since in MSGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$. Similarly, Fig. 4(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3. Fig. 4(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1. For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML. As explained in Section III, our conjecture asserts that any value of $\gamma$ surpassing this lower bound guarantees that MSGT performance will be at least on par with that of ML. Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound. This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection IV-2 provide empirical support for our conjecture. Fig.6(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT. It can be observed that the LVA step performs an extensive filtering process, which allows MSGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
@@
The converse of the GT problem with general prior statistics was developed by Li et al. [52] and according to which any GT algorithm with a maximum error probability $P_{e}$ requires a number of tests that satisfies: $T\leq\left(1-P_{e}\right)H(\mathbf{U})$, where $H(\cdot)$ denotes entropy. Using the joint entropy identity we have: $\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right)$.
@@
The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$. Substituting those priors, it follows that the converse of our problem is: $T\leq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right)$. This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
IV-2 Algorithm Evaluation

### v3

# Non-Adaptive Multi-Stage Algorithm and Bounds for Group Testing with Prior Statistics

@@ Abstract
In this paper, we propose an efficient multi-stage algorithm for non-adaptive Group Testing (GT) with general correlated prior statistics. The proposed solution can be applied to any correlated statistical prior represented in trellis, e.g., finite state machines and Markov processes. We introduce a variation of List Viterbi Algorithm (LVA) to enable accurate recovery using much fewer tests than objectives, which efficiently gains from the correlated prior statistics structure. We also provide a sufficiency bound to the number of pooled tests required by any Maximum A Posteriori (MAP) decoder with an arbitrary correlation between infected items. Our numerical results demonstrate that the proposed Multi-Stage GT (MSGT) algorithm can obtain the optimal MAP performance with feasible complexity in practical regimes, such as with COVID-19 and sparse signal recovery applications, and reduce in the scenarios tested the number of pooled tests by at least $25\%$ compared to existing classical low complexity GT algorithms. Moreover, we analytically characterize the complexity of the proposed MSGT algorithm that guarantees its efficiency.
@@ I Introduction
In this work, we introduce a practical non-adaptive Multi-Stage GT (MSGT) algorithm for correlated items with prior statistics. The proposed multi-stage algorithm employs a new variation of the parallel List Viterbi Algorithm (LVA) [30, 31, 32] we designed for GT to enable accurate low complexity recovery using fewer tests. The proposed algorithm can be applied for any statistical prior represented in trellis [33], e.g., finite-state machines and Markov processes. Using LVA, MSGT leverages those statistics to estimate the defective set efficiently, even in a regime below the maximum likelihood (ML) upper bound. Furthermore, we show how the algorithm’s parameters can be tuned to achieve a maximum probability of success without exceeding the limitation of the available computational capacity. We derive a lower bound that considers the exact priors of the problem and provides analytical results that characterize the MSGT computational complexity efficiency. We also provide an analytical sufficiency bound to the number of pooled tests required by any maximum a-posteriori (MAP) decoder. This bound applies to any arbitrary underlying correlation between infected items. Our numerical results demonstrate that in practical regimes for COVID-19 [34, 35] and sparse signal recovery in signal processing [9, 10, 11, 13], the low computational complexity MSGT algorithm proposed herein can reduce in the scenarios tested the number of pool tests by at least $25\%$.
The rest of this paper is organized as follows. Section II formally describes the GT model with correlated prior statistics. Section III presents the MSGT algorithm and the analytical results, and Section IV presents a sufficiency bound for MAP decoder. Section V details the simulation evaluation. Finally, Section VI provides concluding remarks and future directions.
@@ III Main Results
In this section, we introduce the efficient multi-stage recovery algorithm for any statistical prior represented in a trellis diagram [33], detailed in Algorithm 1. In the first stage, standard low-complexity algorithms [19] reduce the search space independently of prior correlations. This reduction is guaranteed by new analytical results we derive. In the second stage, the algorithm employs a novel adaptation of the List Viterbi Algorithm (LVA) [32], designed for GT to enable accurate low-complexity recovery using fewer tests by exploiting the correlated prior information. Additionally, we derive a bound to ensure the low complexity of the entire algorithm. Section III-A describes the proposed algorithm. Section III-B provides analytical results, followed by a discussion in Section  III-C. For a detailed explanation of all the algorithms used as integral components of Algorithm 1, see Appendix A.
@@ III-B Analytical Results
Subsequently, the probability of success satisfies
@@
The proposition below provides a sufficient condition on $\gamma$, such that, MSGT achieves performance that outperforms ML. This can be assumed since LVA used for this step is an optimal MAP estimator [32].
@@
Let $\eta$ denote the average success probability of ML decoder. For GT with correlated prior information and $\gamma\geq 1$, MSGT achieve an average success probability equal or greater than $\eta$, if
@@
Consider a group test for a population of $N$ items, of which $K$ are defective and a Bernoulli testing matrix. The computational complexity of the MSGT algorithm is bounded by $\mathcal{O}\left(L\gamma^{K}KN\log_{2}N\right)$ operations.
@@ III-C Discussion
IV MAP Analytical Bound for GT with General Correlated Prior Statistics
In this section, we provide a sufficiency bound for the number of required pooled tests for any MAP decoder with an arbitrary underlying correlation between infected items. This bound also applies to the LVA-based MSGT algorithm, which can approach MAP decoder performance for a sufficiently large list [32]. We calculate this bound by extending the known result of maximum likelihood (ML) decoding for Group Testing (GT) problems [17], which bounds the probability of error of a ML decoder using a Gallager-type bound [44] for multiple access channel (MAC) channels [50]. Recall that a noiseless GT setup can be considered as a noiseless MAC setting with $N$ “users”, each having a single codeword. While these type of bounds are well explored in the literature for MAC channels with correlated prior statistics [51, 52, 53, 54, 55, 56, 57], they do not expand naturally to GT, since GT can be thought as a special case where only “$K$” users transmit at any given point. This key difference necessitates dedicated analysis as follows in this section. In particular, to gain from the correlated prior statistics information, unlike ML decoders and bounds e.g., in [17, 14]. We first introduce new notations and definitions in Subsection IV-A, then provide the probability of error to bound the MAP performance in Subsection IV-B, characterizing in Subsection IV-C the maximal number of tests required to guarantee reliability as given in Theorem 5.
IV-A Definitions and notations
Let $\left[n\right]=\left\{1,\ldots,n\right\}$. Let $H(\cdot)$ and $I(\cdot;\cdot)$ denote Shannon’s entropy and mutual information, respectively [58]. All vectors are column vectors and the $\log$ operator denotes $\log_{2}$, unless stated otherwise. Let $\omega\in\left[\binom{N}{K}\right]$ denote the index of a specific set of $K$ items out of the $N$ items. Throughout the paper, $\omega^{*}$ denotes the index of the set of the $K$ actual infected items. Let $W$ denote the random variable which points to the index of the infected items, and has a probability distribution of $P_{W}\left(\cdot\right):\left[\binom{N}{K}\right]\rightarrow\left[0,1\right]$. In the following, $W$ might be omitted for convenience of read. The $\omega$-th set is denoted by $S_{\omega}$. Note that $\lvert S_{\omega}\rvert=K$. The rows of $\mathbf{X}$ that correspond to the items in $S_{\omega}$ are denoted by $\mathbf{X}_{S_{\omega}}$. The output For a given $\omega\neq\omega^{*}$, the true positives (TP)/ false negatives (FN)/ false positives (FP) are denoted by $TP\left(\omega,\omega^{*}\right),FN\left(\omega,\omega^{*}\right),FP\left(\omega,\omega^{*}\right)$. Note that $\lvert FN\left(\omega,\omega^{\prime}\right)\rvert=\lvert FP\left(\omega,\omega^{\prime}\right)\rvert$. In the following, we omit the dependency on $\omega,\omega^{*}$ and write $\mathbf{X}_{TP},\mathbf{X}_{FP},\mathbf{X}_{FN}$ instead of $\mathbf{X}_{TP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FP\left(\omega,\omega^{*}\right)},\mathbf{X}_{FN\left(\omega,\omega^{*}\right)}$ for notational simplicity. We say that an error of size $i$ happens if an output set of items $\omega\neq\omega^{*}$ is chosen and $\lvert FP\left(\omega,\omega^{*}\right)\rvert=i$. This event is denoted by $E_{i}$. Assuming the event $E_{i}$ happened, the index of the FN items is denoted by $j$, i.e. $j\in\left[\binom{N-K}{i}\right]$. Let $\Omega$ denote a function that converts a $j\in\left[\binom{N-K}{i}\right]$ index to a set of size $i$ of the FN items (the dependence on $\omega^{*}$ is omitted for notational simplicity). A decoder is denoted by $g$ and outputs some index $\omega=g\left(Y^{T}\right)$, where the dependency on $\mathbf{X}$ is omitted in favor of notational simplicity. These definitions are illustrated in the following example.
Example
Consider a case with $N=7$ items, where $K=3$ are defective. Suppose the first $3$ items are defective, i.e. items $\left\{1,2,3\right\}$ are defective. In this case, $S_{\omega^{\star}}=\left\{1,2,3\right\}$. Suppose there exists another set that is more likely: $S_{\omega^{\prime}}=\left\{1,4,5\right\}$. In this case, $TP\left(\omega^{\prime},\omega^{*}\right)=\left\{1\right\},FP\left(\omega^{\prime},\omega^{*}\right)=\left\{4,5\right\},\Omega_{2}\left(j\right)=FN\left(\omega^{\prime},\omega^{*}\right)=\left\{2,3\right\}$, where $j$ is the index of the FN set. In this example, an error of size $i=\lvert FP\left(\omega^{\prime},\omega^{*}\right)\rvert=2$ occurred, i.e. the event $E_{2}$ happened.
IV-B Probability of Error
The optimal MAP decoding rule that minimizes the decoding error probability selects an index $\omega$ that satisfies
$P_{W}\left(S_{\omega}\right)P(Y^{T}|\mathbf{X}_{S_{\omega}})\geq P_{W}\left(S_{\omega^{\prime}}\right)P\left(Y^{T}|\mathbf{X}_{S_{\omega^{\prime}}}\right);\hskip 5.69046pt\forall\omega^{\prime}\neq\omega.$ (8)
Recall that the pooling matrix $\mathbf{X}$ is chosen according to a Bernoulli distribution with parameter $p$ in an i.i.d. fashion. An error occurs if there exists $\omega^{\prime}\neq\omega^{*}$ such that $P_{W}\left(\omega^{\prime}\right)P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\prime}}}\right)\geq P_{W}\left(\omega^{*}\right)P\left(Y^{T}\lvert\mathbf{X}_{S_{\omega^{\star}}}\right)$. This event is denoted by $\zeta_{\omega^{\prime}}$. The average probability of error, in terms of exact recovery as defined in Section II, is defined as
$P_{e}=\sum_{\omega}P_{W}\left(S_{\omega}\right)P\left(g\left(Y^{T}\right)\neq\omega\lvert\omega\right).$ (9)
Let $E_{i}^{j}$ denote the event where an error of size $i$ at index $j\in\left[\binom{N}{i}\right]$ happens. The error probability for this event is
$P\left(E_{i}^{j}\right)=\sum_{\omega\in\left[\binom{N}{K}\right]}P\left(w\right)P\left(E_{i}^{j}\lvert w\right).$ (10)
By plugging in (10) into (9), we get
$P_{e}$ $=\sum_{i=1}^{K}\sum_{j=1}^{\binom{N-K}{i}}P\left(E_{i}^{j}\right)$
$=\sum_{i=1}^{K}\sum_{j=1}^{\binom{N-K}{i}}\sum_{\omega=1}^{\binom{N}{K}}P_{W}\left(\omega\right)P\left(E_{i}^{j}\lvert W=\omega\right).$
IV-C Sufficiency GT MAP Bound
Let $(S_{2},S_{1})$ denote a partition of the defective set $S$ into disjoint sets $S_{2}$ and $S_{1}$, with cardinalities $i$ and $K-i$, respectively. This partition allows for categorizing error occurrences into distinct groups, where in group $i$, $K-i$ defective items are already identified correctly, and the primary error scenario arises from trying to detect the remaining $i$ defective items.
Theorem 5.
Consider a group test with a MAP decoder for the population with correlated prior statistics. For $K=O(1)$, if the number of the tests satisfies
$T\geq(1+\varepsilon)\max_{i=1,\ldots,K}\frac{K}{i}H\left(P_{S_{2}|S_{1}}\right)$
then, as $N\rightarrow\infty$ the average error probability approaches zero.
The sufficiency proof of Theorem 5 is twofold. We first bound in the following lemma the error of probability of the optimal MAP decoder as defined in (8). We then use this error of probability analysis to bound the required number of tests required by any decoder.
Lemma 1.
Consider a group test with a MAP decoder for a population with correlated prior statistics. The error probability $P(E_{i}^{j})$ in (10) is bounded by
$P(E_{i}^{j})\leq 2^{-T\left(E_{0}(\rho)-\frac{\frac{1}{\binom{K}{i}}E_{s,j}(\rho,P_{W})}{T}\right)},$
where the error exponent $E_{0}(\rho)$ is given by
$\begin{split}&E_{0}(\rho)=\\ &-\log\sum_{Y^{T}}\sum_{X_{TP}}\left(\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y^{T},X_{TP}\lvert X_{FP}\right)^{\frac{1}{1+\rho}}\right)^{1+\rho},\end{split}$
the correlated prior statistics function of the items $E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)$ is given by
$\begin{split}E_{s,j}\left(\rho,P_{W}\left(\cdot\right)\right)=\log\sum_{S_{1}\in\mathcal{T}\mathcal{P}_{i}}{\left[\sum_{S_{2}\in\mathcal{F}\mathcal{P}_{i}\left(S_{1}\right)}P_{W}\left(\omega^{*}\right)^{\frac{1}{\left(1+\rho\right)}}\right]^{1+\rho}},\end{split}$
for $0\leq\rho\leq 1$.
Proof of Lemma 1: The proof is given in Appendix B. ∎
Proof of Theorm 5.
Define
$\emph{f}(\rho)=E_{o}(\rho)-\frac{E_{s,j}(\rho,P_{W})}{T}.$
Since $0\leq\rho\leq 1$ can be optimized, we aim to show that $Tf(\rho)\to\infty$ as $N\to\infty$ for some $\rho$ within this range. If this holds for all $E_{i}^{j}$, then, due to the resulting exponential decay of $P(E_{i}^{j})$, applying a simple union bound will demonstrate that the total error probability remains small, thus completing the proof of Theorem 5. Since the function $f(\rho)$ is differentiable and admits a power series expansion, we use a Taylor series expansion in the neighborhood of $\rho=0$, yielding
$f(\rho)=f(0)+\rho f^{\prime}(0)+\frac{\rho^{2}}{2}f^{\prime\prime}(\psi)$
for some $\psi\in(0,\rho)$.
For completeness, we show as in [17] that
$\frac{\partial E_{o}}{\partial\rho}|_{\rho=0}$
$=$ $\sum_{Y}\sum_{X_{TP}}[\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\log P\left(Y,X_{TP}|X_{FP}\right)$
-- $\sum_{X_{FP}}P\left(X_{FP}\right)P\left(Y,X_{TP}|X_{FP}\right)\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})]$
$=$ $\sum_{Y}\sum_{X_{TP}}\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})$
$\log\frac{p(Y,X_{TP}|X_{FP})}{\sum_{X_{FP}}P(X_{FP})p(Y,X_{TP}|X_{FP})}=I(X_{FP};X_{TP},Y).$
And also, in Appedix C, that
$\frac{\partial E_{s,j}(\rho,P_{W})}{\partial\rho}|_{\rho=0}=H(P_{S_{2}|S_{1}}).$
This is obtained directly from [59] as also elaborated in [57, Chap. 1]. Note that $E_{o}(0)=0$ and $E_{s,j}(0,P_{W})=0$, hence we have
$T\emph{f}(\rho)=$ $T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)-\frac{H\left(P_{S_{2}|S_{1}}\right)}{T}\Bigg{)}$
-- $\log\frac{1}{\binom{K}{i}}+T\frac{\rho^{2}}{2}E_{o}^{\prime\prime}\left(\psi\right)-T\frac{\rho^{2}}{2}E_{s,j}^{\prime\prime}\left(\psi\right),$
and, if
$T\geq\left(1+\varepsilon\right)\max_{i=1,\ldots,K}\frac{H\left(P_{S_{2}|S_{1}}\right)}{I\left(X_{FP};X_{TP},Y\right)}$ (11)
for some $\varepsilon>0$, we have
$T\emph{f}(\rho)\geq-\log\frac{1}{\binom{K}{i}}$
$+T\rho\Bigg{(}I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)+\frac{\rho}{2}\left(E_{o}^{\prime\prime}(\psi)-E_{s,j}^{\prime\prime}\left(\psi\right)\right)\Bigg{)}$
Moreovver, note that $E_{o}^{\prime\prime}\left(\psi\right)$ is negative [44]. However, it is independent of the other constants and $T$, hence choosing
$0<\rho<\frac{2I\left(X_{FP};X_{TP},Y\right)\left(\frac{\varepsilon}{1+\varepsilon}\right)}{|E_{o}^{\prime\prime}\left(\psi\right)-E_{s,j}^{\prime\prime}\left(\psi\right)|},$
and as $\frac{1}{\binom{K}{i}}$ is fixed and independent of $T$, we have $T\emph{f}(\rho)\rightarrow\infty$ as $N\rightarrow\infty$.
In practice, the expression $I(X_{FP};X_{TP},Y)$ in (11) is important to understand how many tests are required, yet it is not a function of the problem parameters and is bounded in [17] and [45] to get a better handle on $T$.
Claim 1 ([17] and [45]).
For large $K$, and under a fixed input distribution for the testing matrix $\left(\frac{\ln(2)}{K},1-\frac{\ln(2)}{K}\right)$, the mutual information between $X_{FP}$ and $(X_{TP},Y)$ is bounded by $I(X_{FP};X_{TP},Y)\geq\frac{i}{K}$.
Hence, by applying Clain 1 in Eq. (11), we achieves the sufficiency bound on $T$ as provided in Theorem 5,
$(1+\varepsilon)\max_{i=1,\ldots,K}\frac{K}{i}H\left(P_{S_{2}|S_{1}}\right),$
such that, as $N\rightarrow\infty$ the average error probability approaches zero, and reliability is established.
V Numerical Evaluation
@@
This section assesses the performance of the proposed MSGT algorithm by numerical study. First, in Subsection V-1, we provide a numerical evaluation to support our theoretical results and bounds given in Section III. Then, in Subsection V-2, we contrast the performance of MSGT with the one of DD, ML, and MAP in a practical regime of $N$ and $K$. To generate the correlated prior information between adjacent items, we use the Gilbert-Elliot (GE) model [60] (also considered very recently analytically in [61]), which is a stationary Markov chain of order 1 with two states: one representing an error phase and the other an error-free phase. The GE model is characterized by initial probabilities assigned to these two states, denoted as $\pi_{i}\in[0,1]^{2}$, as well as transition probabilities between them $\mathbf{\Phi}_{i}\in[0,1]^{2\times 2}$. These characteristics align well with the inputs required by the MSGT algorithm. In the practical scenarios tested (e.g., in the regime of COVID-19, when the test machine can simultaneously process a fixed small number of measurements [34, 35], or in sparse signal recovery in signal-processing with fixed vector size of input samples [9, 10, 11, 13, 62]), we show that the low computational complexity MSGT algorithm can reduce the number of pool tests by at least $25\%$.
V-1 Theoretical Analysis
@@
In Fig. 4(a) we show the concentration of $\left|\mathcal{P}^{(S_{1,1})}\right|$, as obtained from the simulation, along with the bound on its expectation and on the worst-case that were calculated in Theorems 1 and 2 respectively. Note that the worst-case scenario regarding MSGT is when LVA filters the correct set of defective items. That may happen if the number of possibly defective items exceeds the threshold $\gamma K$. Since in MSGT, we only allow this deviation from the average and ignore the case of exceeding this threshold, our upper bound for the worst-case does not cover all potential realizations of $\left|\mathcal{P}^{(S_{1,1})}\right|$. Similarly, Fig. 4(b) demonstrates the concentration of $\left|\mathcal{P}^{(DD)}\right|$, as acquired through simulation, and the lower bound on its expectation as given in Theorem 3. Fig. 4(c) illustrates the numerically computed lower bound for $\gamma$, derived from the inequality provided in Proposition 1. For this simulation, we fix $N=500$, $K=3$, and calculate $\gamma$ value for a specific range of $T$ relative to the upper bound of ML. As explained in Section III, our conjecture asserts that any value of $\gamma$ surpassing this lower bound guarantees that MSGT performance will be at least on par with that of ML. Therefore, whenever computational resources allow, it is advisable to choose the value of $\gamma$ corresponding to the lower bound. This approach was followed in the subsequent simulations, and the practical outcomes presented in Subsection V-2 provide empirical support for our conjecture. Fig.6(a) compares the number of potential combinations to be examined in the MAP step, with and without the execution of the LVA step in MSGT. It can be observed that the LVA step performs an extensive filtering process, which allows MSGT to remain feasible even when executing MAP is no longer possible, especially in a regime below ML upper bound.
@@
The converse of the GT problem with general prior statistics was developed by Li et al. [63] and according to which any GT algorithm with a maximum error probability $P_{e}$ requires a number of tests that satisfies
$T\geq\left(1-P_{e}\right)H(\mathbf{U}).$
Using the joint entropy identity we have
$\left(1-P_{e}\right)H(\mathbf{U})=\sum_{i=1}^{N}H\left(U_{i}\lvert U_{1},\ldots,U_{i-1}\right).$
@@
The GE model considered in our numerical evaluations is a stationary Markov chain with $\tau=1$. Thus, $P\left(U_{i}\lvert U_{0},\ldots,U_{i-1}\right)=P\left(U_{i}\lvert U_{i-1}\right)$, for $i\in\{2,\ldots,N\}$. Substituting those priors, it follows that the converse of our problem is
$T\geq H(U_{1})+\sum_{i=2}^{N}H\left(U_{i}\lvert U_{i-1}\right).$
This bound is illustrated in the practical scenarios tested in Fig. 4 and Fig. 5.
V-2 Algorithm Evaluation